1 moles ok c atom with gram what is the mass of 1 atom of carbon​...

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Answer:

The value of the mole is equal to the number of atoms in exactly 12 grams of pure carbon-12. 12.00 g C-12 = 1 mole C-12 atoms = 6.022 × 1023 atoms • The number of particles in 1 mole is called Avogadro's Number (6.0221421 x 1023).

Thanks